The gross domestic product (GDP) of a country was $112 billion in 2001. If the GDP grows a constant rate of 4.3% per year. In what year will the GDP be greater than 200 billion.
Answer by lwsshak3(11628)

The gross domestic product (GDP) of a country was $112 billion in 2001. If the GDP grows a constant rate of 4.3% per year. In what year will the GDP be greater than 200 billion.
**
Formula for continuous growth rate: A=Pe^rt, P=initial amt, r=annual growth rate, A=amt after t years.
200=1122e^.043t
200/112=e^.043t
take log of both sides
ln(200/112)=.043t*lne
lne=1
t=ln(200/112)/.043≈13.5
GDP be greater than 200 billion in about 14 years